# Fast and Accurate Decoder for the XZZX code Using Simulated Annealing

### v1

@@ Abstract
The XZZX code is a variant of the surface code designed to address biased noise in realistic quantum devices. For the XZZX code, we propose a decoder based on simulated annealing (SA). Our SA decoder can be readily and efficiently parallelized, by virtue of its simple MCMC-based algorithm. To prepare an initial configuration of SA, we propose to employ recovery chains obtained by a decoder which utilizes a kind of greedy matching graph algorithm. Although $Z$-biased noise is commonly assumed in real quantum devices, we focus on $Y$-biased noise, for which the minimum-weight perfect matching (MWPM) algorithm fails to decode accurately. Our numerical simulation for the code capacity noise model, where only data qubits suffer errors, confirmed that our SA decoder is more accurate than the MWPM decoder. Furthermore, our SA decoder attained the accuracy equivalent to that of the optimal decoder formulated by integer programming, called CPLEX decoder. In our greedy matching decoder, we randomly determine order of matching pairs of incorrect syndromes that have the same distance. This randomness brings about a variety of initial configurations of SA, which leads to faster convergence of our SA decoder. By comparing decoding times of our SA decoder, the CPLEX decoder, and matrix product state (MPS) decoder, all of which can handle $Y$-biased noise appropriately, we confirmed that our SA decoder is fastest if parallelized. This result implies a potential for combining of our greedy matching and SA decoder for practical use in quantum computing.
XZZX code, simulated annealing, biased noise, code capacity noise, maximum likelihood decoder, Metropolis algorithm
@@ I Introduction
The ultimate goal of quantum computing is to realize fault-tolerant quantum computation (FTQC) with robust error correction mechanisms. For superconducting devices, decoherence time (qubit lifetime) is typically on the order of hundreds of microseconds. For this reason, the decoding process must be completed within this time frame. Practical decoders require a simple algorithm to be readily implemented in hardware.
@@
The surface code is one of the most prospective topological error correction codes [1, 2]. It is experimentally feasible, because it requires measuring only nearest-neighbor data qubits [3, 4, 5, 6, 7].
@@
In physical systems such as superconducting qubits, optics and ion traps, realistic noise is biased, that is, $Z$ (phase-flip) errors occur more frequently than $X$ (bit-flip) errors. For such a biased noise, replacing the stabilizer generators of the surface code leads to alternative codes, such as the tailored surface code (which has only $Y$-type stabilizers) [8] and the XZZX code [9]. The XZZX code exhibits higher thresholds for $X$-,$Y$- or $Z$-biased noises, compared with the ordinary surface code, which has stabilizer generators consisting of either only $X$ or $Z$ operators, and is known as the CSS (Calderbank–Shor–Steane) surface code [10, 11]. For the XZZX code, there appeared proposals on its rotated version [12, 13], its cluster state [14] and its physical realization in optical and superconducting devices [15] and neutral atom devices [16].
The minimum-weight perfect matching (MWPM) decoder [1] can be executed in polynomial time of the number of incorrect syndromes and, hence, is well used for the surface code. However, it deals with $X$ and $Z$ errors independently, and is thus less accurate under noises that include $Y$ errors.
On the other hand, MCMC (Markov chain Monte Carlo) decoder and matrix product state (MPS) decoder [17], known as MLD (maximum likelihood decoder), possess high decoding accuracy by taking into account the probability of $Y$ errors. For the $Z$-biased and $Y$-biased noises, the MPS decoder has been researched [18]. Unfortunately, the MPS decoder has a high computational complexity of order $\mathcal{O}(d^{2})$ for the code distance $d$. In contrast, the MCMC decoders have a possibility for fast decoding, since the MCMC algorithms use heuristic approach based on pseudo-random numbers.
@@
One of the earliest studies [19] on MCMC decoders for the CSS surface code employed parallel tempering, while another study [20] utilized the Metropolis method fixed at the Nishimori temperature. For initial configuration of MCMC, they employ error configurations generated by acting stabilizers and logical operators randomly to actual error configuration [19, 12] and recovery chain obtained by the MWPM decoder [20]. Thus, the aim of these previous works is to establish an accurate threshold for the code.
@@
To efficiently minimize energy in MCMC methods, simulated annealing (SA) [21] is often used in statistical physics. By formulating the decoding problem as a random-bond Ising model (RBIM), SA has been applied to decoders for the CSS surface code [22] and the color code [23]. Another formulation of SA decoder for the CSS surface code has hamiltonian with penalty terms [24], and is demonstrated in hardware-implemented annealing machine [25, 26, 22].
@@
The XZZX code has been studied by leveraging the MWPM decoder, the MPS decoder [9] and MCMC decoder [12]. They investigated the decoding accuracy for $Z$-biased noise, which is typical in real quantum devices. Assuming knowing the probabilities of $X,Y,Z$ errors, they incorporate this prior information into weight coefficients of the hamiltonian so as to improve decoding accuracy.
@@
The research [12] assumes the code capacity noise model, in which only data qubits suffer errors independently and identically. To explore configurations efficiently, they carried out the MCMC method at lower inverse temperature than the Nishimori temperature, keeping the weight coefficients invariant in the hamiltonian. Their numerical experiment suggests that calculating a dominant term (composed of degeneracy and energy) of the partition function is sufficient to determine the correct logical operator class. Thus, their algorithm may be useful to collect this term to utilize it as train data in neural network decoder.
@@
In the present article, we propose a simulated annealing (SA) decoder for the XZZX code. The aim of our research is not to determine an accurate threshold, but rather to establish a decoding algorithm that is moderately accurate and yet sufficiently fast for practical use on a quantum computer. We assume the code capacity noise. Specifically, our interest is the case that each qubit experiences $Y$-biased noise, which cannot be decoded appropriately by the MWPM decoder. As initial configuration for SA, we propose to employ a recovery chain obtained from a decoder which utilizes a greedy matching algorithm. We implement the whole algorithms in the C++ language from scratch, in order to attain decoding time as short as possible.
@@
The rest of the present paper is organized as follows. In Sec. II, we review the XZZX code. For the XZZX code, we describe principle of error detection and correction in Sec. III and decoding algorithms, including our SA decoders in Sec. IV. We show numerical result and argue decoding accuracy and time in Sec. V and discuss the benefits of our SA decoder in Sec. VI. In Sec. VII we conclude.
@@ II The XZZX code
In this section, we briefly review the XZZX surface code [9]. We consider a rotated square lattice with open boundaries [27]. In the lattice, data qubits reside on the vertices.
@@
Figure 1: The XZZX code for the code distance $d=5$. (a) Stabilizer generators $G_{f}$, showing four-body in the bulk and three-body operators at the boundaries. (b) Logical operator $L_{Z}$ (c) Logical operator $L_{X}.$
@@
whose form is the origin of the name ‘XZZX code’. This stabilizer generator is a four-body operator in bulk and a three-body operator at the boundaries. The XZZX code has only one type of stabilizer generators, unlike the CSS surface code which has two types of generators. Since each of the stabilizer generator includes both $X$ and $Z$ operators as shown in Eq. (1), the XZZX code is not CSS-type. The number of stabilizer generators is equal to the number of faces, i.e., $N_{\mathrm{stabilizer}}=2d(d-1)=\mathcal{O}(d^{2})$.
At the center of every face $f$, an ancilla qubit is located to detect errors. For each ancilla qubit, we measure the stabilizer generator $G_{f}$.
We define logical operators $L_{Z}$ ($L_{X}$) as a chain connecting the left and right (upper and lower) boundaries, respectively. Fig. 1(1),(1) depict examples of the logical operators $L_{Z}$ and $L_{X}$. Furthermore, we define logical operator $L_{Y}\coloneqq\mathrm{i}\,L_{Z}\,L_{X}$. As long as the opposite boundaries are connected, we regard these chains as logical operators, namely, allow the freedom of a stabilizer operator in the stabilizer group $\mathcal{G}\coloneqq\langle G_{f}\,|\,f:\text{face}\rangle$. More precisely, for $P\in\{I,X,Z,Y\}$, the logical operator is defined as equivalence class $L_{P}]$ by imposing invariance under the action of the stabilizer group $\mathcal{G}$ in the sense that $G\,L_{P}]=[L_{P}]$ for any $G\in\mathcal{G}$.
The code distance $d$ is defined as the length of the shortest representation of the logical operators $L_{Z}$ and $L_{X}$. For the surface codes including the XZZX code, it is the linear length of the lattice. The number of data qubits is $N_{\mathrm{q}}=d^{2}+(d-1)^{2}$.
The XZZX code is advantageous under biased noises. That is, the XZZX code has threshold $p_{\mathrm{th}}=0.5$ for pure $X$, $Y$, $Z$ noises, because it is equivalent to classical repetition codes; On the other hand, the CSS surface code has $p_{\mathrm{th}}=0.109$ for the pure $X,Z$ noises [1] and $p_{\mathrm{th}}=0.5$ for the pure $Y$ noise [8].
III Principle of Error detection and correction
Throughout the present paper, error configuration means the information that each data qubit suffers which $I$ (no error), $X$, $Z$ or $Y$ error. Fig. 2 illustrates an example of the error configuration. In Fig. 2, X,Y,ZX,\,Y,\,Z errors occur on specific data qubits.
@@ III Principle of Error detection and correction
The vertices are the centers of faces with the color $\bullet$ and are the ancillas for the stabilizer measurement.
@@
The edges which connect nearby vertices in $\mathcal{V}_{\bullet}$. The edge centers represent errors which suffer data qubits.
@@
The set of edges which suffer errors is called error chain. There are two types of error chains $C_{\mathrm{white}}\subset\mathcal{E}_{\mathrm{white}}$ and $C_{\mathrm{gray}}\subset\mathcal{E}_{\mathrm{gray}}$, which connect centers of white and gray face, respectively, drawn with the pink and purple lines in Fig. 2. For both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the horizontal (vertical) edges represent $Z$ ($X$) errors, respectively. Solely on the intersection points of the horizontal and vertical edges in the different types of error chains, $Y$ errors appear.
@@
Figure 2: An example of the error configuration. The circles indicate qubits which suffer errors with error type $X$,$Y$,$Z$. (The qubits with no errors are not drawn.) The error configuration is expressed as two types of error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, connecting the white faces and gray faces and drawn with the pink and purple lines, respectively. The endpoints of these error chains are captured by the incorrect syndromes, drawn with the green squares. In both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the $Z$ ($X$) errors reside on horizontal (vertical) edges, respectively. The $Y$ error appears solely on the intersection point of both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$.
@@
For each face $f$, we measure the stabilizer generator $G_{f}$ and thereby obtain a measurement outcome $s_{f}\in\{0,1\}$. For an error chain $C_{\bullet}$, it holds the eigenvalue equation
[…]
@@
In Fig. 2, incorrect syndromes are drawn with green squares. The incorrect syndromes capture the endpoints of $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$. In the endpoint of horizontal (vertical) edges, $Z$ ($X$) errors anti-commute with $X$ ($Z$) operators acting on the left and right (up and down) qubits in the stabilizer generator $G_{f}$ and hence can be detected.
@@
For an error chain $C_{\bullet}$, the set of face centers whose ancillas possess incorrect syndromes is called syndrome and denoted by
$S_{\bullet}\coloneqq\left\{(\text{center of face $f$})\in\mathcal{V}_{\bullet}\,|\,s_{f}=1\right\}.$ (3)
The syndrome consistency can then be expressed as
@@
The two error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$ contain the same information as the error configuration CC [^1] In the present article, to distinguish an error configuration $C$, we denote an error chain by $C_{\bullet}$ with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$. . We express the correspondence as $C=(C_{\mathrm{white}},C_{\mathrm{gray}})$ and then define the tuple $\partial{C}\coloneqq(\partial{C_{\mathrm{white}}},\partial{C_{\mathrm{gray}}})$. Similarly, we define the tuple $S\coloneqq(S_{\mathrm{white}},S_{\mathrm{gray}})$. Using this abbreviated notation, the syndrome consistency of Eq. (4) reads $\partial{C}=S$.
The actual error configuration $C$ is unknown; The only information available is the syndrome $S$, from which decoders infer a recovery configuration $C$ such that $\partial{C}=S$. Decoders can be classified into two types: those that identify the maximum likelihood error (MLE) configuration and those that identify the maximum-likelihood logical operator class. The latter is referred to as the maximum likelihood decoder (MLD). The examples of the MLE decoder are the MWPM decoder (Sec. IV.2) and our greedy matching decoder (Sec. IV.3). The examples of MLD are our SA decoders (Sec. IV.1, Sec. IV.4), decoder formulated as integer programming (Sec. A) and the MPS decoder [17]. In general, the second type is more accurate than the first type, because it can take into account the degeneracy of error configurations with the same energy.

### v2

@@ Abstract
The XZZX code is a variant of the surface code tailored to address biased noise in realistic quantum devices. We propose a simulated annealing (SA) decoder for the XZZX code. Our SA decoder is amenable to parallelization because its MCMC updates are simple and local. To initialize SA, we use a recovery configuration produced by our greedy matching decoder. Although $Z$-biased noise is commonly assumed in realistic quantum devices, we instead focus on $Y$-biased noise, where MWPM becomes suboptimal because it neglects correlations induced by $Y$ errors. Our numerical simulations for the code capacity noise model, where only data qubits suffer errors, show that our SA decoder achieves higher accuracy than the MWPM decoder. Furthermore, our SA decoder achieves an accuracy comparable to that of the optimal minimum-energy (MAP-configuration) decoder formulated as an integer programming problem, called the CPLEX decoder. In our greedy matching decoder, we randomize the tie-breaking among equal-weight pairs. This randomness generates a variety of initial configurations for SA, which can lead to faster convergence of our SA decoder. By comparing decoding times of our SA decoder, the CPLEX decoder, and the matrix product state (MPS) decoder, all of which can handle $Y$-biased noise appropriately, we estimate that our SA decoder could be competitive in runtime under an idealized assumption of near-perfect parallel efficiency. These results suggest that combining SA with our greedy matching initializer is a practical approach toward fault-tolerant quantum computation.
XZZX code, simulated annealing, biased noise, code capacity noise model, Metropolis algorithm
@@ I Introduction
The ultimate goal of quantum computing is to realize fault-tolerant quantum computation (FTQC) with robust error correction mechanisms. In superconducting qubit platforms, coherence time (i.e., qubit lifetime) is typically on the order of hundreds of microseconds. The error correction process must be completed within this time frame. Practical decoders therefore favor simple algorithms to be readily implemented in hardware.
@@
The surface code is one of the most promising topological error correction codes [kitaev, surface_code_towards]. It is experimentally feasible because it requires only nearest-neighbor interactions between data and ancilla qubits [Zhao, Suppressing_experiment, iOlius, google_surface_code, 17qubit_experiment].
@@
In physical systems such as superconducting qubits, optical systems, and ion traps, realistic noise is often biased, meaning that $Z$ (phase-flip) errors occur more frequently than $X$ (bit-flip) errors. For such biased noise, modifying the stabilizer generators of the surface code leads to alternative codes, such as the tailored surface code (which has only $Y$-type stabilizers) [tuckett_tailoring] and the XZZX code [xzzx_nature]. The XZZX code achieves higher thresholds under $X$-, $Y$-, or $Z$-biased Pauli noise than the conventional surface code. The conventional surface code is a CSS (Calderbank–Shor–Steane) code [calderbank_shor, steane] with stabilizer generators consisting solely of $X$-type or $Z$-type operators. For the XZZX code, there have been proposals for its rotated version [xzzx_agnostic, rotated_xzzx], a cluster-state (foliated) construction [foliated_xzzx], numerical simulations for optical and superconducting qubits [Darmawan_xzzx] and neutral-atom qubits [neutral_atom].
The minimum-weight perfect matching (MWPM) decoder [kitaev] can be executed in polynomial time in the number of syndrome defects and, hence, is widely used for the surface code. However, in the standard MWPM approach, since it does not capture the correlations induced by $Y$ errors, its performance can degrade when $Y$ errors are non-negligible.
On the other hand, an MCMC (Markov chain Monte Carlo) decoder and a matrix product state (MPS) decoder [efficient_MLD] exhibit high decoding accuracy by taking into account the probability of $Y$ errors. For $Z$- and $Y$-biased noise, the MPS decoder has been studied [ultrahigh]. However, for a fixed truncation (bond) dimension $\chi$, the MPS decoder scales as $\mathcal{O}(d^{2})$ (up to polynomial factors in $\chi$). In contrast, MCMC-based decoders can be fast in practice because each MCMC update is local and simple, and the computation is easily parallelized.
@@
One of the earliest studies [Wootton] on MCMC decoders for the CSS surface code employed parallel tempering, while another study [Hutter] utilized the Metropolis method fixed at the Nishimori inverse temperature. For an initial error configuration of MCMC, they employed error configurations generated by acting stabilizers and logical operators randomly on the true error configuration [Wootton, xzzx_agnostic] and a recovery chain obtained by the MWPM decoder [Hutter]. Thus, the aim of these previous works was to determine the threshold of the code accurately.
@@
To efficiently minimize energy in MCMC methods, simulated annealing (SA) [Kirkpatrick] is often used in statistical physics. By formulating the decoding problem as a random-bond Ising model (RBIM), SA has been applied to decoders for the CSS surface code [takeuchi_comparative_decoder] and the color code [takada]. Another formulation of an SA decoder for the CSS surface code has a Hamiltonian with penalty terms [fujii_penalty], and was demonstrated in a hardware-implemented annealing machine [fujisaki_DA, fujisaki_circuit-level, takeuchi_comparative_decoder].
@@
The XZZX code has been studied using the MWPM decoder, the MPS decoder [xzzx_nature], and the MCMC-based decoders [xzzx_agnostic]. They investigated the decoding accuracy for $Z$-biased noise, which is typical in quantum hardware platforms. Assuming knowledge of the probabilities of $X,Y,Z$ errors, the authors incorporated this prior information into the weight coefficients of the Hamiltonian so as to improve decoding accuracy.
@@
Ref. [xzzx_agnostic] assumes the code capacity noise model, in which only data qubits suffer independently and identically distributed errors. To explore configurations efficiently, they carried out the MCMC method at a lower inverse temperature than the Nishimori inverse temperature while keeping the weight coefficients invariant in the Hamiltonian. Their numerical results suggest that extracting a dominant term of the partition function is sufficient to determine the correct logical operator class. Thus, their algorithm may be useful for collecting this term, which could then be used as training data for a neural-network decoder.
@@
In this paper, we propose a simulated annealing (SA) decoder for the XZZX code. Our goal is not to determine the threshold as accurately as possible, but rather to develop a decoding algorithm that is reasonably accurate, yet sufficiently fast for practical use on a quantum computer. We assume the code capacity noise model. Specifically, we are interested in the case where each qubit experiences $Y$-biased noise, a regime in which the MWPM decoder becomes significantly suboptimal because it neglects correlations induced by $Y$ errors. To initialize SA, we employ a recovery configuration produced by a greedy matching decoder. All algorithms are implemented in C++ to minimize decoding time.
@@
The rest of this paper is organized as follows. In Sec. II, we review the XZZX code. For the XZZX code, we describe the principles of error detection and correction in Sec. III, and then present decoding algorithms which include our SA decoders in Sec. IV. We show numerical results in Sec. V and discuss the benefits of our SA decoder in Sec. VI. We conclude in Sec. VII.
@@ II The XZZX code
In this section, we briefly review the XZZX surface code [xzzx_nature]. We consider a rotated square lattice with open boundaries [kitaev-open_boundary]. Data qubits reside on the vertices.
@@
Figure 1: The XZZX code for the code distance $d=5$: (a) Stabilizer generators $G_{f}$, showing four-body operators in the bulk and three-body operators at the boundaries. (b) Logical operator $L_{Z}$. (c) Logical operator $L_{X}$.
@@
whose form is the origin of the name “XZZX code”. This stabilizer generator is a four-body operator in the bulk and a three-body operator at the boundaries. The XZZX code has only one type of stabilizer generators, unlike the CSS surface code which has two types of generators. Since each stabilizer generator includes both $X$ and $Z$ operators as shown in Eq. (1), the XZZX code is not a CSS code. At the center of every face $f$, an ancilla qubit is placed to measure the stabilizer generator $G_{f}$ and detect errors.
We define logical operators $L_{Z}$ ($L_{X}$) as a Pauli chain connecting the left and right (top and bottom) boundaries, respectively. Fig. 1(1) and (1) depict examples of the logical operators $L_{Z}$ and $L_{X}$, respectively. We further define $L_{Y}\coloneqq\mathrm{i}\,L_{Z}\,L_{X}$.
Throughout this paper, we ignore global phases of Pauli operators (i.e., overall factors $\pm 1$ and $\pm\mathrm{i}$ are identified). Let $\overline{\mathcal{P}}$ denote the $N_{\mathrm{q}}$-qubit Pauli group modulo global phases. In particular, we identify $L_{Y}$ with $L_{Z}L_{X}$ in $\overline{\mathcal{P}}$.
We identify Pauli chains that differ by multiplication of a stabilizer operator in the stabilizer group $\mathcal{G}\coloneqq\langle G_{f}\mid f:\text{face}\rangle$. More precisely, for $P\in\{I,X,Y,Z\}$, we define the logical Pauli operator as the equivalence class $L_{P}]$ modulo stabilizers, i.e., $L_{P}\sim GL_{P}$ for any $G\in\mathcal{G}$.
The code distance $d$ is defined as the minimum weight of a nontrivial logical operator (equivalently, the length of the shortest Pauli chain representing $L_{Z}$ or $L_{X}$). For the surface codes including the XZZX code, this distance coincides with the linear size of the lattice. The number of data qubits is $N_{\mathrm{q}}=d^{2}+(d-1)^{2}$. The number of stabilizer generators is equal to the number of faces, i.e., $N_{\mathrm{stabilizer}}=2d(d-1)=\mathcal{O}(d^{2})$.
The XZZX code is advantageous under biased noise. That is, the XZZX code has a threshold $p_{\mathrm{th}}=0.5$ for pure $X$, $Y$, or $Z$ noise in the code capacity noise model, reflecting its mapping to a classical repetition code [xzzx_nature]; on the other hand, the CSS surface code has $p_{\mathrm{th}}=0.109$ for the pure $X$ or $Z$ noise [kitaev] and $p_{\mathrm{th}}=0.5$ for the pure $Y$ noise [tuckett_tailoring].
III Principles of error detection and correction
Throughout this paper, an error configuration $C\in\overline{\mathcal{P}}$ specifies, for each data qubit, whether it experiences no error ($I$) or a Pauli error $X$, $Y$, or $Z$. Fig. 2 illustrates an example of the error configuration. In Fig. 2, X,Y,ZX,\,Y,\,Z errors occur on specific data qubits.
@@ III Principles of error detection and correction
The vertices are the centers of faces with the color $\bullet$; each vertex corresponds to an ancilla qubit used to measure the stabilizer.
@@
The edges connect nearest-neighbor vertices in $\mathcal{V}_{\bullet}$. Each edge is associated with the data qubit located at its midpoint; an error on that data qubit is represented on the corresponding edge.
@@
The set of edges on which errors occur is called an error chain. There are two types of error chains $C_{\mathrm{white}}\subset\mathcal{E}_{\mathrm{white}}$ and $C_{\mathrm{gray}}\subset\mathcal{E}_{\mathrm{gray}}$, which connect centers of white and gray faces, respectively, and are drawn with the pink and purple lines in Fig. 2. In this paper, to avoid confusion, $C$ (without a subscript) denotes an error configuration, whereas $C_{\bullet}$ (with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$) denotes an error chain. For both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the horizontal and vertical edges represent the presence of $Z$- and $X$-type Pauli components, respectively (i.e., $Z$ or $Y$ on a horizontal edge, and $X$ or $Y$ on a vertical edge). A $Y$ error occurs exactly at the intersection point of both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$. With a slight abuse of notation, we use the same symbol $C_{\bullet}$ to denote the corresponding Pauli operator.
@@
Figure 2: An example of the error configuration $C$. Circles indicate data qubits with $X$, $Y$, or $Z$ errors; qubits without errors are omitted. The error configuration can be represented as two types of error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, connecting the white faces and gray faces and drawn with the pink and purple lines on the corresponding decoding graphs $(\mathcal{V}_{\mathrm{white}},\mathcal{E}_{\mathrm{white}})$ and $(\mathcal{V}_{\mathrm{gray}},\mathcal{E}_{\mathrm{gray}})$, respectively. The endpoints of these error chains are captured by the syndrome defects, drawn with the green squares. In both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the $Z$- and $X$-type Pauli components correspond to operators on data qubits located at the centers of horizontal and vertical edges, respectively. A $Y$ error occurs precisely on a data qubit where both the $Z$- and $X$-type components are present (i.e., where the corresponding horizontal and vertical edges are both occupied).
@@
For each face $f$, we measure the stabilizer generator $G_{f}$ and thereby obtain a measurement outcome $s_{f}\in\{0,1\}$. For an error chain (viewed as a Pauli operator) $C_{\bullet}$, the outcome is determined by whether $C_{\bullet}$ commutes or anti-commutes with $G_{f}$:
$G_{f}\,C_{\bullet}=(-1)^{s_{f}}\,C_{\bullet}\,G_{f},$ (2)
i.e., $s_{f}=0$ if $G_{f},C_{\bullet}]=0$ and $s_{f}=1$ if $\{G_{f},C_{\bullet}\}=0$. We refer to a face $f$ with $s_{f}=1$ as a syndrome defect.
A syndrome defect appears precisely at an endpoint of the error chain $C_{\bullet}$. More precisely, $s_{f}=1$ if and only if $v_{f}\in\partial C_{\bullet}$, where $v_{f}$ denotes the vertex (i.e., face center) corresponding to the face $f$, and $\partial C_{\bullet}$ denotes the set of vertices in $\mathcal{V}_{\bullet}$ that have odd incidence with the edge set $C_{\bullet}$ (equivalently, the endpoints of the chain $C_{\bullet}$).
@@
In Fig. 2, syndrome defects are drawn with green squares. They capture the endpoints of $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$. At such endpoints of horizontal and vertical edges, $Z$- and $X$-type error components anti-commute with $X$ and $Z$ operators acting on the left/right and up/down qubits in the stabilizer generator $G_{f}$, respectively, and hence can be detected.
@@
For an error chain $C_{\bullet}$, the set of face centers whose ancilla qubits possess syndrome defects is called the syndrome and denoted by
$S_{\bullet}\coloneqq\left\{v_{f}\in\mathcal{V}_{\bullet}\,\middle|\,s_{f}=1\right\}.$ (3)
The syndrome-consistency condition can then be expressed as
@@
The two error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$ contain the same information as the error configuration $C$. We express the correspondence as $C=(C_{\mathrm{white}},C_{\mathrm{gray}})$ and then define the tuple $\partial{C}\coloneqq(\partial{C_{\mathrm{white}}},\partial{C_{\mathrm{gray}}})$. Similarly, we define the tuple $S\coloneqq(S_{\mathrm{white}},S_{\mathrm{gray}})$. Using this abbreviated notation, the syndrome-consistency condition of Eq. (4) reads $\partial{C}=S$.
The true error configuration $\bar{C}$ is unknown; the only available information is the syndrome $S$, from which decoders infer a recovery configuration $C$ such that $\partial{C}=S$.